# Strassen Attention: Unlocking Compositional Abilities in Transformers Based on a New Lower Bound Method

### v2

@@ A.1.1 Function Composition
  • $n$ is sampled uniformly from the range $[N_{\text{min}},N_{\text{max}}]$, with $N_{\text{min}}=25$ and $N_{\text{max}}=30$.
@@
A.1.2 Binary Relation Composition
@@
  • $X=\textsf{flatten}(R)$ is an input sequence of length $n=m^{2}$ of a flatten boolean matrix $R\in\{0,1\}^{m\times m}$ representing the binary relation. Each element $R_{ij}$ is independently set to $1$ with a probability $P$ between 00 to $1$, where $P$ is a parameter independent on the input length.
  • $m$ is sampled uniformly from the range $[N_{\text{min}},N_{\text{max}}]$, with $N_{\text{min}}=6$ and $N_{\text{max}}=8$.
@@
The dataset generation process ensures randomness in $X$ while adhering to the constraints for $Y$. For $m$ in the range $[N_{\text{min}},N_{\text{max}}]$, the probability $P$ is set to $32.5\%$ to achieve a balanced proportion of positive labels. Algorithm 2 provides the detailed data generation procedure.
@@ A.1.3 Match3
  • $X=(x_{0},x_{1},\dots,x_{n-1})$ is a pseudo-random sequence of integers sampled from the range $[0,M-1]$, where $M=37$.
@@
  • $n$ is sampled uniformly from the range $[N_{\text{min}},N_{\text{max}}]$, where $N_{\text{min}}=30$ and $N_{\text{max}}=35$.
The dataset generation process aims to balance the distribution of ones in $Y$ across four predefined bins corresponding to percentage ranges: $[0,25)\%$, $[25,50)\%$, $[50,75)\%$, and $[75,100]\%$. Algorithm 3 outlines the data generation procedure.
@@
  Initialize four empty bins {Each bin corresponding to percentage ranges of ones in sequences: $[0,25)\%$, $[25,50)\%$, $[50,75)\%$, and $[75,100]\%$}
@@
  • The input sequence $X=(6,9,9,9,7,10,9,34,9,9,30,\dots)$ contains pseudo-random integers between 00 and $M-1$ ($M=37$).
@@
A.1.4 Quotient Binary Relation Composition
@@
  • col is a function that assigns a unique color to each element in $[0,m-1]$, with colors sampled uniformly from the range $[0,m-1]$.
  • $m$ is sampled uniformly from the range $[N_{\text{min}},N_{\text{max}}]$, with $N_{\text{min}}=6$ and $N_{\text{max}}=8$.
@@
The dataset generation process ensures randomness in $R$ while adhering to the constraints for $Y$. For $m$ in the range $[N_{\text{min}},N_{\text{max}}]$, the probability $P$ is set to $43.3\%$ to achieve a balanced proportion of positive labels. Algorithm 4 provides the detailed data generation procedure.
@@
     for each pair $(i,j)$ in $[0,m-1]\times[0,m-1]$ do
        Append $1$ to $Y$ if $i\neq j$ and there exist $k_{1},k_{2}$ such that:
@@
        Otherwise, if $i\neq j$, append 00 to $Y$, else append $-100$ to $Y$
@@ A.2 Experimental Protocol
We implement all models using PyTorch framework [1] with Opt-Einsum library [5]. We employ AdamW optimizer for all training tasks without a learning rate scheduler, ensuring a consistent optimization strategy across experiments. We use the Binary Cross-Entropy loss as the objective function. In order to handle padding, we used attention masks, preventing the model from attending to padded positions, and thus ensuring no changes in the model’s outputs. For the target sequence, we assign a value of $-100$ to positions corresponding to padded tokens. This ensures that during loss and accuracy calculations, only tokens with values other than $-100$ are considered. We achieve this by applying a mask, where predictions and targets are filtered as pred = pred[mask] and target = target[mask], respectively.
@@ Appendix B Further Experiments
  • Standard and Triangular ($n=$Nodes×$\times$×Nodes) GPU memory usage does not shows a bottleneck even for configurations with hidden dimensions of 2048 and input lengths of 16,384.
  • Triangular ($n=$Nodes) Limitations appear for hidden dimensions of 1024 or higher and input lengths as low as 196, with increasing severity for longer input sequences.

### v3

@@ 7.1.1 Function Composition
  • $n$ is sampled uniformly from the range $N_{\text{min}},N_{\text{max}}]$, with $N_{\text{min}}=25$ and $N_{\text{max}}=30$.
@@
7.1.2 Binary Relation Composition
@@
  • $X=\textsf{flatten}(R)$ is an input sequence of length $n=m^{2}$ of a flatten boolean matrix $R\in\{0,1\}^{m\times m}$ representing the binary relation. Each element $R_{ij}$ is independently set to $1$ with a probability $P$ between 0 to $1$, where $P$ is a parameter independent on the input length.
  • $m$ is sampled uniformly from the range $N_{\text{min}},N_{\text{max}}]$, with $N_{\text{min}}=6$ and $N_{\text{max}}=8$.
@@
The dataset generation process ensures randomness in $X$ while adhering to the constraints for $Y$. For $m$ in the range $N_{\text{min}},N_{\text{max}}]$, the probability $P$ is set to $32.5\%$ to achieve a balanced proportion of positive labels. Algorithm 2 provides the detailed data generation procedure.
@@ 7.1.3 Match3
  • $X=(x_{0},x_{1},\dots,x_{n-1})$ is a pseudo-random sequence of integers sampled from the range [0,M1][0,M-1], where $M=37$.
@@
  • $n$ is sampled uniformly from the range $N_{\text{min}},N_{\text{max}}]$, where $N_{\text{min}}=30$ and $N_{\text{max}}=35$.
The dataset generation process aims to balance the distribution of ones in $Y$ across four predefined bins corresponding to percentage ranges: $0,25)\%$, $25,50)\%$, $50,75)\%$, and $75,100]\%$. Algorithm 3 outlines the data generation procedure.
@@
 Initialize four empty bins {Each bin corresponding to percentage ranges of ones in sequences: $0,25)\%$, $25,50)\%$, $50,75)\%$, and $75,100]\%$}
@@
  • The input sequence $X=(6,9,9,9,7,10,9,34,9,9,30,\dots)$ contains pseudo-random integers between 0 and M1M-1 ($M=37$).
@@
7.1.4 Quotient Binary Relation Composition
@@
  • col is a function that assigns a unique color to each element in [0,m1][0,m-1], with colors sampled uniformly from the range [0,m1][0,m-1].
  • $m$ is sampled uniformly from the range $N_{\text{min}},N_{\text{max}}]$, with $N_{\text{min}}=6$ and $N_{\text{max}}=8$.
@@
The dataset generation process ensures randomness in $R$ while adhering to the constraints for $Y$. For $m$ in the range $N_{\text{min}},N_{\text{max}}]$, the probability $P$ is set to $43.3\%$ to achieve a balanced proportion of positive labels. Algorithm 4 provides the detailed data generation procedure.
@@
  for each pair $(i,j)$ in $0,m-1]\times[0,m-1]$ do
   Append 11 to $Y$ if $i\neq j$ and there exist $k_{1},k_{2}$ such that:
@@
   Otherwise, if $i\neq j$, append 0 to $Y$, else append 100-100 to $Y$
@@ 7.2 Experimental Protocol
We implement all models using PyTorch framework [1] with Opt-Einsum library [5]. We employ AdamW optimizer for all training tasks without a learning rate scheduler, ensuring a consistent optimization strategy across experiments. We use the Binary Cross-Entropy loss as the objective function. In order to handle padding, we used attention masks, preventing the model from attending to padded positions, and thus ensuring no changes in the model’s outputs. For the target sequence, we assign a value of 100-100 to positions corresponding to padded tokens. This ensures that during loss and accuracy calculations, only tokens with values other than 100-100 are considered. We achieve this by applying a mask, where predictions and targets are filtered as pred = pred[mask] and target = target[mask], respectively.
@@ 8 Further Experiments
[…]
  • Triangular (n=n=Nodes) Limitations appear for hidden dimensions of 1024 or higher and input lengths as low as 196, with increasing severity for longer input sequences.
9 Future directions and limitations
Our results pave the way for several directions of future research. First, we have introduced a new method to obtain limitations for one-layer Transformer, based on splitting the VC dimension. We hope that the method could be applied to other architectures and other tasks involving complex reasoning in more naturalistic settings (e.g., semantic parsing). Second, given the differences observed in learning stability between the third-order and Strassen attentions, the latter seems to be associated with a smoother loss landscape, an hypothesis that needs to be confirmed and studied. Third, Strassen attention can be adapted to incorporate interactions that involve more than three tokens, possibly capturing more complex patterns in data. Yet, practical learnability of such higher-order interactions needs to be assessed. Finally, and related to the previous point, although the main goal of our work was to gain a deeper theoretical understanding of the abilities of the Transformer, our conclusions are limited by using toy tasks. Our next step is to test the theoretical advantages of the Strassen attention using specific benchmarks or even in learning methods such as masked language modeling. At the same time, the possible applications of the Strassen attention are not limited to compositionality but could extend to such areas as knowledge graphs, protein structure prediction and others.